sin z-tanz Find and classify all singularities of the function f(z) = 2
Answer #1

solution Given that (2) = sine-tana If a function is not analytic at 2=0 then f (2) is said to have singularity at 2 = 0 f(a)122222222 Cost = 0 7 = don+ MEI OT 7 = (49+1) A , MEI te (antu MEI for different values of n, it will give different & therefcase 2 = 0 This will be a type of isolated singularity as the function will not be analytic at only t=0 Now ex = 1 - 4-0 Now